OpenStudy (hy123):

Find the general solution of each of the following differential equations. If an initial condition is given, find the particular solution that satisfies this initial condition. y' + 2ty = 1, y(0)=1

OpenStudy (hy123):

@thomaster @RadEn

OpenStudy (abdela25):

you want to find the value of y ??? if not could you please explain your question more because i cannot understand it very good

OpenStudy (hy123):

Well it is a differential equation and i need to solve for y

OpenStudy (hy123):

@ganeshie8

ganeshie8 (ganeshie8):

linear differential equation

ganeshie8 (ganeshie8):

start by finding the integrating factor

OpenStudy (hy123):

Yes.

OpenStudy (hy123):

Wouldn't that be. y' = 1 - 2ty?

OpenStudy (hy123):

Making it a standard form and then diving both sides by y?

ganeshie8 (ganeshie8):

y' + p(x) y = q(x) integrating factor = \(\large e^{\int p ~dx}\)

OpenStudy (hy123):

hm.. okay.

ganeshie8 (ganeshie8):

y' + 2ty = 1 integrating factor = \(\large e^{\int 2t ~dt} = e^{t^2}\)

ganeshie8 (ganeshie8):

multiply that both sides

ganeshie8 (ganeshie8):

\(\large y' + 2ty = 1 \) multiply both sides wid the integrating factor \(\large e^{t^2}\) \(\large y'e^{t^2} + ye^{t^2}2t = e^{t^2} \)

OpenStudy (hy123):

Got that.

OpenStudy (hy123):

Then do we take the integral of both sides?

ganeshie8 (ganeshie8):

yes, and observe that left side is simply a derivative of product : \(ye^{t^2}\)

ganeshie8 (ganeshie8):

\(\large y' + 2ty = 1 \) multiply both sides wid the integrating factor \(\large e^{t^2}\) \(\large y'e^{t^2} + ye^{t^2}2t = e^{t^2} \) \(\large (ye^{t^2})' = e^{t^2} \) integrate both sides now..

OpenStudy (hy123):

The integration of e^t^2 is a bit hard.

OpenStudy (hy123):

Do you mind showing me how you got it.

ganeshie8 (ganeshie8):

nope, its simply a new funciton

OpenStudy (hy123):

oh

OpenStudy (hy123):

e^t^2 F(t) + C?

ganeshie8 (ganeshie8):

\(\large \int e^{x^2}dx = \frac{\sqrt{\pi}}{2} erfi(x)\)

OpenStudy (hy123):

what is erfi?

ganeshie8 (ganeshie8):

\(\large \int e^{x^2}dx = \frac{\sqrt{\pi}}{2} erfi(x) + C \)

ganeshie8 (ganeshie8):

its a new function like logx and many other functions defined to be the area under the bell curve

OpenStudy (hy123):

Ok.

ganeshie8 (ganeshie8):

\(\large y' + 2ty = 1 \) multiply both sides wid the integrating factor \(\large e^{t^2}\) \(\large y'e^{t^2} + ye^{t^2}2t = e^{t^2} \) \(\large (ye^{t^2})' = e^{t^2} \) integrate both sides now.. \(\large \int (ye^{t^2})' = \int e^{t^2} \) \(\large ye^{t^2} = \frac{\sqrt{\pi}}{2} erfi(t) + C\)

OpenStudy (hy123):

Then you divide both side by e^t^2.

ganeshie8 (ganeshie8):

before that, find the value of C, apply initial conditions...

OpenStudy (hy123):

Plugging in y(0) = 1 and solving for C?

ganeshie8 (ganeshie8):

yes

OpenStudy (hy123):

I get C = -1/2

ganeshie8 (ganeshie8):

how ?

OpenStudy (hy123):

Actually lemme re-do it.

ganeshie8 (ganeshie8):

okay

OpenStudy (hy123):

Actually i got 1...

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

plug that in and solve y

OpenStudy (hy123):

C = 1? Correct?

ganeshie8 (ganeshie8):

correct

OpenStudy (hy123):

Do you mind showing me what you got?

OpenStudy (hy123):

Nvm i got it. Now i got to hed to class. Thanks a lot Ganeshie.

ganeshie8 (ganeshie8):

\(\large y' + 2ty = 1 \) multiply both sides wid the integrating factor \(\large e^{t^2}\) \(\large y'e^{t^2} + ye^{t^2}2t = e^{t^2} \) \(\large (ye^{t^2})' = e^{t^2} \) integrate both sides now.. \(\large \int (ye^{t^2})' = \int e^{t^2} \) \(\large ye^{t^2} = \frac{\sqrt{\pi}}{2} erfi(t) + C\) ***** y(0) = 1 \(\large 1e^{0} = \frac{\sqrt{\pi}}{2} erfi(0) + C\) \(\large 1 = 0 + C\) \(\large 1 = C\) \(\large ye^{t^2} = \frac{\sqrt{\pi}}{2} erfi(t) + 1\) \(\large y = e^{-t^2} ( \frac{\sqrt{\pi}}{2} erfi(t) + 1)\)

ganeshie8 (ganeshie8):

u wlc :)